Please show all work
NAME (print neatly): Score (last name, first name) Markup Problem. Consider a situation, where a machine shop is bidding a contract for to provide an aluminum manifold. The costs per unit are $51.36 and the price the component will sell for is $96.42. Important to show all work and provide appropriate units with two decimal places. 1. What is the markup based on cost and price? (1 point for each part) Markup as a percentage of cost Markup as a percentage of price 2. How many units must be sold to achieve a profit of $175k? (2 points) 3. Now, assume costs of $57.51, what would the unit price need to be if the manufacturer used a markup as a percentage of the price of 39%? (3 points) Now, assume a price of $91.77, what would the unit costs need to be if the manufacturer used a markup as a percentage of the cost of 39%13 points) 4.
Answer #1
answer 1)
Markup as percentage of cost = =(96.42-51.36)/51.36
Markup as percentage of cost = 87.73%
Markup as percentage of Price = =(96.42-51.36)/96.42
Markup as percentage of Price = 46.73%
answer 2)
Profit per unit = =(96.42-51.36) 45.06
required profit = 175000
Number of unit = required profit/ profit per unit                               3,884
answer 3)
Cost = 57.51
Mark up on price = 39%
required price = 57.51/(1-39%)                               94.28
answer 4)
Price per unit = 91.77
Mark up as percentage on cost = 39%
unit cost = =91.77/(1+39%)                               66.02